OpenStudy (anonymous):

Q#1: 2x + 5x = -9 + 4 Q #2: 4y 6y --- = ---- 2 3 Q#3: 32x + 2 = 4x +3

OpenStudy (callisto):

Solve the equations?

OpenStudy (anonymous):

Help :( This is new to me.

OpenStudy (callisto):

Do you know how to do addition for like terms?

OpenStudy (anonymous):

no, just help :9 plz plz plz

OpenStudy (callisto):

Example: x+2x = 3x Now, you can try this: : 2x + 5x = ..?

OpenStudy (anonymous):

x = 7?!

OpenStudy (callisto):

Not really. x + 2x = 3x ^ You just ''cover'' the x and do addition for 1 and 2, so you get 3. Then add an ''x'' after the number 3. So, when you add x and 2x together, you get 3x 2x+5x is on the left, you need to add them together first. You also have the right side of the equation to simplify...

OpenStudy (anonymous):

oh ok , so please hep with those questions! 2x + 5x = -9 + 4 4y 6y --- = ---- 2 3 32x + 2 = 4x +3 ANSWER THEM.

OpenStudy (callisto):

I'm already helping you....

OpenStudy (anonymous):

do the questions

OpenStudy (callisto):

No free answers.... For the first question, do additions for the left and right, then divide both sides by the coefficient of x.

OpenStudy (anonymous):

4y 6y --- = ---- 2 3 For this one , am i right? : 4y x 3 = 6y x 2 12y = 12y y = 12 (6, 4, 3, 2) ---- 12 (6, 4, 3,2) y = 2?

OpenStudy (callisto):

No.. You can't solve it.. since it is 12y=12y...

OpenStudy (anonymous):

or i think its y = 1

OpenStudy (anonymous):

Knew it! Then, please help me dear friend :)

OpenStudy (callisto):

y can be any numbers... since 12y = 12y => y=y (of course y is y :\)

OpenStudy (anonymous):

ok so was i right if "y can be any numbers'' ?!

OpenStudy (anonymous):

???

OpenStudy (callisto):

That's not the way to say it :\

OpenStudy (anonymous):

Ok , fine sorry. Please tell me! This is important :(

OpenStudy (callisto):

It's ''infinite solutions'', at least it's the way I answer it...

OpenStudy (anonymous):

2x + 5x = -9 + 4 4y 6y --- = ---- 2 3 32x + 2 = 4x +3

OpenStudy (callisto):

Have you tried the first one?

OpenStudy (anonymous):

the 2x + 5x = -9 + 4?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

No! This is new to me. If you would be nice and do all the questions for me then it would gladly make my day? Thanks

OpenStudy (anonymous):

Thank you! <3

OpenStudy (anonymous):

i got it . And sorry callisto..

OpenStudy (anonymous):

you deleted it?

OpenStudy (anonymous):

32x + 2 = 4x +3 for this one you do 32x - 4x because you subtract 4x on both sides to get if over to the side with 32x then you subtract 2 on both sides then solve for x like i did in the last one

OpenStudy (anonymous):

where is the first one you answered

OpenStudy (anonymous):

so 32x - 4x = 28x?

OpenStudy (anonymous):

i didnt delete it this site gave me a "warning" and deleted it

OpenStudy (callisto):

''2x + 5x = -9 + 4 you add the 2x and 5x and then the -9 + 4 so it would look like 7x = -5, then you divide -5 by 7 to...'' Credits: @blondie16

OpenStudy (anonymous):

Will you write again?

OpenStudy (anonymous):

thanks @Callisto

OpenStudy (anonymous):

Thanks Callisto :) Sorry!

OpenStudy (anonymous):

Will it be -5/7 first or 7/-5?

OpenStudy (callisto):

Divide a by b = a/b

OpenStudy (anonymous):

so its -5/7?

OpenStudy (anonymous):

but @sellymiranda so 32x - 4x = 28x? is not right... you have the 32x - 4x right but you can put those together to = 28x but your not finished there... 32x + 2 = 4x +3 28x = 1 then divide 1 by 28

OpenStudy (anonymous):

how u got the 1?

OpenStudy (anonymous):

4-3?

OpenStudy (anonymous):

+2 you gotta subtract the 2 from the 3 after you move the 4x over

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its 28/1 or 1/28?

OpenStudy (anonymous):

1/28

OpenStudy (anonymous):

thats 0.03.etc

OpenStudy (anonymous):

Shouldn't it be 28/1?

OpenStudy (anonymous):

you can just leave it as a fraction

OpenStudy (anonymous):

so x = 28

OpenStudy (anonymous):

1/28

OpenStudy (anonymous):

so x = 1/28?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i kind of need help with this; '2x + 5x = -9 + 4 you add the 2x and 5x and then the -9 + 4 so it would look like 7x = -5, then you divide -5 by 7 to...''

OpenStudy (anonymous):

how??